a particle moves along straight line find distance travelled in. time interval o to 5sec.
1.v=t^2-4 

Grade:12th pass

1 Answers

Khimraj
4 years ago
v =  ds/dt = t2 – 4
ds = (t2 – 4)dt
integrating form 0 to 5 sec
s = [t3/3 – 4t]o5 = 125/3 – 20 = 65/3 m